Answers Given Please Show WorkTwo hot-dog shops with constant average and marginal costs c= 0 are located on a street of length 1. Let x = 0 represent the
Answer #1

Ans The demand of q2 is based on location as shown in question. The transport cost is not given and after solving the above question we will get the slope of p2 = 1.667

And the Nash equilibrium price for shop 2 is = 0.31

According to formula p2=

(x2- x1 )/3 (4l - x1- x2)

Where x1& x2 are location which are given in question

X1= 0.3

X2= 0.6

And l is the total of location which is given = 1

So putting the value in the equation we get :

(0.6- 0.3)/3 (4*1-0.3-0.6)

0.3/3 (4-0.9)

1/10(3.1)

= 0.31

If the travel cost down then the price for both shops goes down and profit will also goes down

According to this model if transport cost is 0 then profit is also 0 for both the firm ad if transport cost increasing then profit is also increasing.

The total demand will be :

P1 + (x1 - X )²= p2 + (x2- X)²

X= x1+x2/2 + p2- p1/2(x2- x1)

X= 0.45 +( 0.31 - 0.09)/ 0.6

X = 0.45+ 0.3667

~ 0.8167